I have 5 types of flowers and three flowers will be in each vase. How many vases will there be?

Answer by ewatrrr(24785)

Hi,
.
*Note: The number of permutations of n distinct objects taken r at a time, where repetitions are not allowed is n!/(n-r)!
.
n!/(n-r)!
In this case 5 types of flowers taken 3 at a time
.
5!/2! =( 5*4*3*2*1)/2*1 = 5*4*3 = 60
.
60 vases